given z1 = 2<45 degrees, z2 =3<120 degrees and z3 =4<180 degrees. determine the following and leave your answer in rectangular form.

  1. (z1)2+z2/z2+z3
  2. z1/z2z3


1
Expert's answer
2022-05-26T11:01:46-0400

1.


"(z_1)^2=(2)^2\angle(2\cdot45\degree)=4\angle90\degree=4i"

"z_2+z_3=3\angle120\degree+4\angle180\degree"

"=3(-\dfrac{1}{2}+\dfrac{\sqrt{3}}{2}i)+4(-1)"

"=-\dfrac{11}{2}+\dfrac{3\sqrt{3}}{2}i"


"\dfrac{z_2}{z_2+z_3}=\dfrac{-\dfrac{3}{2}+\dfrac{3\sqrt{3}}{2}i}{-\dfrac{11}{2}+\dfrac{3\sqrt{3}}{2}i}=\dfrac{3-3\sqrt{3}i}{11-3\sqrt{3}i}"

"=\dfrac{(3-3\sqrt{3}i)(11+3\sqrt{3}i)}{121+27}"

"=\dfrac{33+9\sqrt{3}i-33\sqrt{3}i+27}{148}"

"=\dfrac{15}{37}-\dfrac{6}{37}i"


"(z_1)^2+\dfrac{z_2}{z_2+z_3}=4i+\dfrac{15}{37}-\dfrac{6}{37}i"


"=\dfrac{15}{37}+\dfrac{142}{37}i"



2.


"z_2z_3=3\angle120\degree(4\angle180\degree)""=3(4)\angle(120\degree+180\degree)=12\angle300\degree"

"=12(\dfrac{1}{2}-\dfrac{\sqrt{3}}{2}i)=6-6\sqrt{3}i"

"\dfrac{z_1}{z_2z_3}=\dfrac{2}{12}\angle(45\degree-300\degree)"

"=\dfrac{1}{6}(\cos(-255\degree)+i\sin(-255\degree))"

"=\dfrac{1}{6}(\cos(105\degree)+i\sin(105\degree))"

"\dfrac{z_1}{z_2z_3}=\dfrac{\sqrt{2}+\sqrt{2}i}{6-6\sqrt{3}i}"


"=\dfrac{(\sqrt{2}+\sqrt{2}i)(6+6\sqrt{3}i)}{36+108}"

"=\dfrac{(\sqrt{2}+\sqrt{6}i+\sqrt{2}i-\sqrt{6})}{24}"

"=-\dfrac{\sqrt{6}-\sqrt{2}}{24}+\dfrac{\sqrt{2}+\sqrt{6}}{24}i"